40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Sat Mar 25 22:43:46 2017, johannes, Noise Hunting, Cavity, CryoCoatingCavs Ringdowns, Error Signals & Co 6x
    Reply  Tue Apr 4 08:18:31 2017, johannes, Noise Hunting, Noise Budget, LB1005 internal noise LB1005_GainSettings_comp.pdfLB1005_East_noise.pdfLB1005_West_noise.pdfbeat_sense_noise.pdf
       Reply  Wed Apr 5 23:27:12 2017, johannes, Noise Hunting, Noise Budget, Sensing noise error_noise_west_20170405b.pdferror_noise_east_20170405b.pdferror_noises_west_20170405b.pdferror_noises_east_20170405b.pdf
          Reply  Fri Apr 7 01:44:22 2017, johannes, Noise Hunting, Noise Budget, Sensing noise 6x
             Reply  Fri Apr 7 02:09:31 2017, Koji, Noise Hunting, Noise Budget, Sensing noise 
          Reply  Sat Apr 8 12:53:33 2017, johannes, Notes, PDH, RIO planex amplitude response laser_response_20170407.pdf
             Reply  Mon Apr 17 16:27:54 2017, johannes, Notes, PDH, RIO planex amplitude response TFAG4395A_17-04-2017_152730.pdfcurrent_driver_mod.pdf
                Reply  Wed Apr 19 12:48:06 2017, johannes, Notes, Laser, Current driver oscillations current_driver_mod.pdfAG4395A_18-04-2017_181713.pdf
                   Reply  Fri Apr 28 11:37:41 2017, johannes, Notes, Laser, RFPD transfer functions RFPD_response_20170413.pdf
                      Reply  Thu May 4 19:13:04 2017, johannes, Notes, Laser, RIO planex frequency response laser_frequency_response.pdfpll_scheme.pdflaser_response_20170504.pdf
                         Reply  Tue May 9 16:18:02 2017, rana, Notes, Laser, LB1005 Transfer Function 
                            Reply  Thu Jun 15 13:58:17 2017, johannes, Notes, Laser, Laser frequency response west_response_20170614.pdf
                      Reply  Wed Jun 14 18:37:15 2017, johannes, Notes, Laser, RFPD circuitry RFPD_notches.jpgRFPD_WEST.jpgRFPD_EAST.jpg
                         Reply  Thu Jun 15 09:24:06 2017, Chris, Notes, Laser, RFPD circuitry 
                            Reply  Thu Jun 15 11:39:03 2017, Koji, Notes, Laser, RFPD circuitry 
                         Reply  Sat Jun 17 02:56:26 2017, johannes, Notes, Laser, RFPD circuitry IMG_20170616_143847305.jpg
                            Reply  Sun Jul 9 21:56:07 2017, johannes, Notes, Laser, RFPD circuitry resPD_20170622.pdf
Message ID: 1567     Entry time: Wed Apr 19 12:48:06 2017     In reply to: 1566     Reply to this: 1573
Author: johannes 
Type: Notes 
Category: Laser 
Subject: Current driver oscillations 

[Rich, Johannes]

Following the previous observations and Rich's suspicion that the observed resonance comes from an oscillation in the HF modulation, we opened up a spare current driver that wasn't in use (but supposed to be identical to the ones used for the cryo cavities) and found that it didn't come as advertised. The attached schematic shows it's shortcomings. The capacitor C39 had been modified (had a test point attached and capacitance added with a secondary cap.), and R40 was still at 1kOhm as in the original version (Note 1 for version 6 on the document says it has been replaced with 0 Ohm due to high frequency oscillations - duh). We placed 100 pF and 0 Ohm as supposed, which removed the 8MHz feature in the transfer function.

I proceeded to apply this fix to both in-use current drivers, the resulting amplitude modulation transfer function is attached. The >1MHz region is still not feature-less, but the ripples do not exceed 5dB above the DC response, as opposed to 40-50 dB before, and the phase looks smooth.

04/19 edit: Along with the change to the current driver, I also changed the lowpass on the demodulation to the less aggressive mini circuits SLP-21.4+, which achieves stopband attenuation at ~60MHz and is therefore sufficient to filter the 2omega terms. With these changes I was able to bring the UGF of the cavity locks up to 300 kHz.

Attachment 1: current_driver_mod.pdf  78 kB  | Hide | Hide all
current_driver_mod.pdf
Attachment 2: AG4395A_18-04-2017_181713.pdf  26 kB  | Hide | Hide all
AG4395A_18-04-2017_181713.pdf
ELOG V3.1.3-